# Bridging Stereo Geometry and BEV Representation with Reliable Mutual Interaction for Semantic Scene Completion

### v3

@@ Abstract
3D semantic scene completion (SSC) is an ill-posed perception task that requires inferring a dense 3D scene from limited observations. Previous camera-based methods struggle to predict accurate semantic scenes due to inherent geometric ambiguity and incomplete observations. In this paper, we resort to stereo matching technique and bird’s-eye-view (BEV) representation learning to address such issues in SSC. Complementary to each other, stereo matching mitigates geometric ambiguity with epipolar constraint while BEV representation enhances the hallucination ability for invisible regions with global semantic context. However, due to the inherent representation gap between stereo geometry and BEV features, it is non-trivial to bridge them for dense prediction task of SSC. Therefore, we further develop a unified occupancy-based framework dubbed BRGScene, which effectively bridges these two representations with dense 3D volumes for reliable semantic scene completion. Specifically, we design a novel Mutual Interactive Ensemble (MIE) block for pixel-level reliable aggregation of stereo geometry and BEV features. Within the MIE block, a Bi-directional Reliable Interaction (BRI) module, enhanced with confidence re-weighting, is employed to encourage fine-grained interaction through mutual guidance. Besides, a Dual Volume Ensemble (DVE) module is introduced to facilitate complementary aggregation through channel-wise recalibration and multi-group voting. Our method outperforms all published camera-based methods on SemanticKITTI for semantic scene completion.
@@ 3.4 Semantic Scene Completion
$*{}^{*}$ (2017) Stereo-PTS 10.31 0.187
$*{}^{*}$ (2020) Stereo-PTS 10.45 0.214
$*{}^{*}$ (2022) Stereo 12.82 0.274
$*{}^{*}$ (2023) Stereo 13.06 0.313
$*{}^{*}$ (2023) Stereo 13.57 0.338
@@
Table 2: Evaluation results of stereo variants. For $*{}^{*}$ and $*{}^{*}$, We employ stereo images as inputs. For $*{}^{*}$ and $*{}^{*}$, we leverage stereo depth net to generate pseudo point clouds (Stereo-PTS).
@@ 4.2 Performance
Evaluation of Stereo Variants. To ensure fair comparisons, we also implement stereo variants of the baselines as shown in Table 2. For $*{}^{*}$ Cao and de Charette (2022) and $*{}^{*}$ Huang et al. (2023), we employ left and right images to generate stereo-based predictions. Other LiDAR-based baselines including LMSCNet Roldao et al. (2020) and SSCNet Song et al. (2017) require 3D geometric inputs, so we adapt them with pseudo-3D inputs leveraging the same GwcNet Guo et al. (2019) used in our framework. Our proposed BRGScene efficiently outperforms all the other methods. For more details on model complexity, please refer to the Supplementary Material.

### v4

@@ Abstract
3D semantic scene completion (SSC) is an ill-posed perception task that requires inferring a dense 3D scene from limited observations. Previous camera-based methods struggle to predict accurate semantic scenes due to inherent geometric ambiguity and incomplete observations. In this paper, we resort to stereo matching technique and bird’s-eye-view (BEV) representation learning to address such issues in SSC. Complementary to each other, stereo matching mitigates geometric ambiguity with epipolar constraint while BEV representation enhances the hallucination ability for invisible regions with global semantic context. However, due to the inherent representation gap between stereo geometry and BEV features, it is non-trivial to bridge them for dense prediction task of SSC. Therefore, we further develop a unified occupancy-based framework dubbed BRGScene, which effectively bridges these two representations with dense 3D volumes for reliable semantic scene completion. Specifically, we design a novel Mutual Interactive Ensemble (MIE) block for pixel-level reliable aggregation of stereo geometry and BEV features. Within the MIE block, a Bi-directional Reliable Interaction (BRI) module, enhanced with confidence re-weighting, is employed to encourage fine-grained interaction through mutual guidance. Besides, a Dual Volume Ensemble (DVE) module is introduced to facilitate complementary aggregation through channel-wise recalibration and multi-group voting. Our method outperforms all published camera-based methods on SemanticKITTI for semantic scene completion. Our code is available on https://github.com/Arlo0o/StereoScene.
@@ 3.4 Semantic Scene Completion
SSCNet (2017) Stereo-PTS 10.31 0.187
LMSCNet (2020) Stereo-PTS 10.45 0.214
MonoScene (2022) Stereo 12.82 0.274
TPVFormer (2023) Stereo 13.06 0.313
OccFormer (2023) Stereo 13.57 0.338
@@
Table 2: Evaluation results of stereo variants on the SemanticKITTI validation set. For MonoScene and TPVFormer, We employ stereo images as inputs. For SSCNet and LMSCNet, we leverage stereo depth net to generate pseudo point clouds (Stereo-PTS).
@@ 4.2 Performance
Evaluation of Stereo Variants. To ensure fair comparisons, we also implement stereo variants of the baselines as shown in Table 2. For MonoScene Cao and de Charette (2022) and TPVFormer Huang et al. (2023), we employ left and right images to generate stereo-based predictions. Other LiDAR-based baselines including LMSCNet Roldao et al. (2020) and SSCNet Song et al. (2017) require 3D geometric inputs, so we adapt them with pseudo-3D inputs leveraging the same GwcNet Guo et al. (2019) used in our framework. Our proposed BRGScene efficiently outperforms all the other methods. For more details on model complexity, please refer to the Supplementary Material.
@@ 5 Conclusion
6 Acknowledgements
This paper is supported in part by NSFC under Grant 62302246 and ZJNSFC under Grant LQ23F010008.
Supplementary Material for BRGScene
[…]
Figure 7: More visualization results on the SemanticKITTI validation set. The overlay shadow areas at the bottom of semantic predictions denote unseen scenery out of the camera’s field of view (FOV). From top to below demonstrate 3 cases. For each one, the first row illustrates stereo input images while the second row depicts comparison results of our BRGScene and Voxformer-S.
Figure 8: Implementation details of the stereo representation construction. The left and right features are encoded with camera parameters before forming the disparity volume.
Appendix A More Visualization Results
Below we provide qualitative comparisons on the SemanticKITTI validation set and show visualization examples in Figure 7. Compared to VoxFormer-SLi et al. (2023a), our BRGScene generates more precise predictions (e.g. cars in example 1; crossroads in example 2 and example 3) and more complete scenes with proper hallucinations (e.g. shadow regions in example 2 and example 3).
Specifically, in example 1, compared to VoxFormer-S, our method predicts the more accurate shapes for cars. In example 2, compared to VoxFormer-S, our method successfully hallucinates a more complete and reasonable scene outside the camera FOV. In example 3, our BRGScene precisely captures the semantics and geometry of the crossroad and the terrain while VoxFormer-S faces challenges in correct perception.
Appendix B Implementation Details on Dual Volume Construction
In this section, we describe implementation details for our dual volume construction. The Stereo Geometric Volume Constructor targets to match 2D extracted features and outputs a regularized depth volume. The BEV Latent Volume Constructor constructs a 3D latent volume coupled with context features from 2D images.
More Details on Stereo Representation Construction. As mentioned in Sec. 3.2 of the main paper, we construct stereo geometric volume from the left and right image features. As shown in Figure 8, the left and right image features are correlated to form the group-wise disparity volume, which is then converted into the depth volume. To regularize the depth volume, we adopt 3D CNNs with hourglass (encoder-decoder) structures to process it Guo et al. (2019); Shen et al. (2021). We first feed the depth volume to two 3$\times$3$\times$3 convolutions with stride 1. Next, we employ three stacked 3D hourglasses to further aggregate the volumetric information. The structure details of the 3D hourglass are shown in Table 5, where two residual connections are adopted to aggregate context from different semantic levels. In each 3D hourglass, the input volume is first downsampled and then upsampled with de-convolution, thus aggregating information along the spatial and depth dimensions. Finally, two 3$\times$3$\times$3 convolutions are used to generate the regularized depth volume.
Name Layer Properties Dim Input
Conv1 $3\times 3\times 3$, 2, 1 $32/64$ Input Volume
Conv2 $3\times 3\times 3$, 1, 1 $/64$ 64 Conv1
Conv3 $3\times 3\times 3$, 2, 1 $64/128$ Conv2
Conv4 $3\times 3\times 3$, 1, 1 $/128$ 128 Conv3
Deconv5 $3\times 3\times 3$, 2, 1 $128/64$ Conv4
Conv6 $1\times 1\times 1$, 1, 0 $/64$ 64 Conv2
Shortcut Addition & ReLU $/64$ 64
Deconv5,
Conv6
Deconv7 $3\times 3\times 3$, 2, 1 $64/32$ Shortcut
Conv8 $1\times 1\times 1$, 1, 0 $/32$ 32 Input Volume
Output Volume Addition & ReLU $/32$ 32
Deconv7,
Conv8
Table 5: Structure details of the 3D hourglass. The """"Layer Properties"""" indicates kernel size, stride and padding, respectively.
Figure 9: Implementation details of the BEV representation construction. We employ two branches in parallel to generate the BEV context features and BEV latent volume.
More Details on BEV Representation Construction. As mentioned in Sec. 3.2 of the main paper, we construct BEV context features and BEV latent volume with the BEV constructor Philion and Fidler (2020); Li et al. (2023b). In our implementation, we adopt left image features and camera parameters as inputs.
As shown in Figure 9, two branches are employed in parallel, where the upper branch is dedicated to generating contextual features while the lower branch is dedicated to generating latent volume.
In the parameter encoder, we concatenate camera intrinsic and extrinsic as camera parameters $P_{i}$. For the purpose of being aware of distinct camera inputs, we feed the camera parameters $P_{i}$ to both branches for camera awareness similar to Li et al. (2023b). Formally,
$P_{e}=\sigma\left(Conv(Reshape(FC(P_{i})))\right)$ (11)
where $Conv$ and $FC$ are convolutions and fully-connected layers, whereas $\sigma$ and $Reshape$ represent sigmoid function and reshape operation, respectively. The encoded camera parameters $P_{e}$ are multiplied with the extracted features to generate the camera-aware features.
In the upper branch, a 3$\times$3 convolution is adopted to generate the context features. In the lower branch, we adopt two stacked residual blocks and an ASPP Chen et al. (2017, 2018) module to refine the depth distribution of the latent volume. The ASPP module aims to expand the receptive field in the depth perception process.
Methods Input mIoU (%) $\uparrow$
MonoScene Mono 11.50
VoxFormer-S Stereo 12.35
VoxFormer-T Stereo-T 13.35
BRGScene (ours) Stereo 15.43
SSCNet LiDAR 16.35
LMSCNet LiDAR 17.19
JS3CNet LiDAR 22.67
Table 6: Comparison with LiDAR-based methods on the SemanticKITTI validation set.
Appendix C Comparison with LiDAR-based methods.
We report the comparison with LiDAR-based methods in Table 6. The table illustrates that the performance gaps between our proposed method and recent LiDAR-based methods are relatively small (e.g., 0.92 mIoU drop compared to SSCNet). These results demonstrate that our proposed stereo-based method can be promoted as a cheap alternative to LiDAR with competitive performance.
Appendix D More Evaluation with Current Structure
We compare other formations of BEV volume in Tab. 7. Although stronger backbones (e.g. BEVFormer and BEVDepth) show slight improvement over LSS, the boosted performance is marginal compared with our proposed complementary volume exploitation structure.
Volume
LSS
BEVFormer
BEVDepth
LSS+Stereo
IoU(%) $\uparrow$ 38.97 39.33 39.42 43.85
mIoU(%) $\uparrow$ 13.18 13.44 13.51 15.43
Table 7: The effect of volume representations with different BEV baselines.